# Med3DVLM: An Efficient Vision-Language Model for 3D Medical Image Analysis

### v1

@@ Abstract
Vision-language models (VLMs) have shown promise in 2D medical image analysis, but extending them to 3D remains challenging due to the high computational demands of volumetric data and the difficulty of aligning 3D spatial features with clinical text. We present Med3DVLM, a 3D VLM designed to address these challenges through three key innovations: (1) DCFormer, an efficient encoder that uses decomposed 3D convolutions to capture fine-grained spatial features at scale; (2) SigLIP, a contrastive learning strategy with pairwise sigmoid loss that improves image-text alignment without relying on large negative batches; and (3) a dual-stream MLP-Mixer projector that fuses low- and high-level image features with text embeddings for richer multi-modal representations. We evaluate our model on the M3D dataset, which includes radiology reports and VQA data for 120,084 3D medical images. Results show that Med3DVLM achieves superior performance across multiple benchmarks. For image-text retrieval, it reaches 61.00% R@1 on 2,000 samples, significantly outperforming the current state-of-the-art M3D-LaMed model (19.10%). For report generation, it achieves a METEOR score of 36.42% (vs. 14.38%). In open-ended visual question answering (VQA), it scores 36.76% METEOR (vs. 33.58%), and in closed-ended VQA, it achieves 79.95% accuracy (vs. 75.78%). These results highlight Med3DVLM’s ability to bridge the gap between 3D imaging and language, enabling scalable, multi-task reasoning across clinical applications. Our code is publicly available at https://github.com/mirthAI/Med3DVLM.
@@ 1 Introduction
Medical image analysis plays a crucial role in diagnosing and treating diseases such as cancer, cardiovascular conditions, and neurological disorders. However, existing models are often task-specific, lack adaptability to new tasks, and do not support real-time user interactions. Vision-language models (VLMs), such as Contrastive Language-Image Pre-Training (CLIP)[27] and Large Language and Vision Assistant (LLaVA)[19], offer greater versatility in dynamic clinical settings by aligning medical images with textual reports. CLIP leverages contrastive learning on image-text pairs, enabling zero-shot image classification and image-text retrieval, while LLaVA extends CLIP by integrating its visual encoder with a large language model (LLM), facilitating interactive tasks such as report generation and visual question answering (VQA). Despite their success in analyzing 2D medical images, such as chest X-rays[8] and pathology slides[21], their application to 3D imaging remains limited.
@@ 2.1 Medical Vision-Language Models
Recent efforts have aimed to develop VLMs for 3D medical imaging. RadFM [38] introduced a generalist VLM trained on massive datasets that consisted of both 2D and 3D medical images and associated radiology reports. It integrated contrastive learning and generative modeling across diverse imaging modalities (X-ray, CT, MRI) and textual reports, enabling a unified radiology representation. Extending this, M3D [2] introduced a large-scale dataset, M3D-Data, along with M3D-LaMed, a multi-modal vision-language model for 3D medical image analysis tasks such as image-text retrieval, report generation, VQA, and promotable segmentation. Its successor, E3D-GPT [14], employed a 3D multi-modal masked autoencoder framework to further enhance image-text retrieval, leading to improved performance in report generation and VQA.
@@ 3.2.1 Contrastive Pretraining
Such a decomposition strategy enables the processing of 3D volumes at a larger size of 128x256×256 in this study, preserving fine-grained spatial details for improved image-text alignment. In this work, we utilize DCFormer-small variant, which consists of a stem stage and four hiararchical stages with [2, 3, 6, 2] layers and [96, 192, 384, 768] channels [1].
@@ 3.2.2 Multi-modal Projector Pretraining
We propose an effective multi-modal projector for the interaction between visual and textual embeddings. In this pretraining stage, all model weights are frozen except for those in the projector. To effectively capture spatial and feature dependencies, we adopt the MLP-Mixer architecture [29], which alternates between a token-mixing MLP (spatial mixer) and a channel-mixing MLP. To the best of our knowledge, this is the first use of the MLP-Mixer in 3D VLMs.
@@
To further enhance multi-modal fusion, we introduce a low-high hybrid Mixer-MLP architecture (Figure 2), inspired by Janus Pro [5]. Unlike M3D-LaMed [2], which produces large spatial outputs $(B,2048,768)$ that require lossy downsampling, DCFormer outputs semantically rich features $(B,32,768)$ from its final layer and spatially detailed features $(B,256,384)$ from its penultimate layer. To leverage both types of features, we utilize two parallel MLP-Mixer modules—one for each layer. After passing through $N$ Mixer layers, the resulting feature outputs are concatenated into a unified sequence of image tokens, which are then combined with text tokens from the tokenizer and fed into the LLM.
This hybrid design enables the projector to fuse high-level abstract semantics with low-level spatial details, improving multi-modal alignment. Unlike prior approaches that downsample features at the expense of context, our method preserves diverse information without increasing token length. As a result, our projector produces richer and more semantically aligned joint representations, significantly enhancing radiology report generation and VQA.
@@ 3.3 Evaluation Metrics
We evaluate the quality of radiology report generation and open-ended VQA using four key metrics: BLEU [24], ROUGE [16], METEOR [3], and BERTScore [43]. BLEU measures precision by evaluating the overlap of n-grams between the generated text and the reference text, focusing on how many words or phrases in the generated report are found in the reference text. ROUGE emphasizes recall, assessing how much of the reference text appears in the generated output, which is particularly useful for summarization tasks to ensure key details are retained. METEOR improves upon BLEU and ROUGE by considering synonyms and stemming, balancing precision and recall to provide a more flexible and comprehensive measure. BERTScore utilizes deep learning-based contextual embeddings to evaluate semantic similarity.
@@ 3.4 Implementation Details
We implemented Med3DVLM using PyTorch [25] and Hugging Face Transformers [37], leveraging DeepSpeed ZeRO2 and BF16 precision on 8 NVIDIA A100 80GB GPUs. We used the AdamW optimizer [20] with gradient accumulation steps of 1, a warmup ratio of 0.03, a cosine learning rate scheduler, and weight decay where specified.
During contrastive pretraining on the M3D-Cap dataset, we resized 3D CT volumes to $128\times 256\times 256$, used DCFormer-small [1] as the vision encoder, and ClinicalBERT [33] as the text encoder. Image features were aggregated via mean pooling, while text features were derived from the [CLS] token, as it captures the global semantic representation of the entire sequence in ClinicalBERT, making it suitable for contrastive alignment with image features. Both were projected to a shared 768-dimensional space. We used a batch size of 64, learning rate of $1\times 10^{-4}$, weight decay of 0.1, and ran for 100 epochs.
For VLM pretraining, we introduced a hybrid MLP-Mixer projector, which fused low-level features from the penultimate encoder layer ($256\times 384$) and high-level features from the final layer ($32\times 768$). Each feature stream passed through $N$ Mixer layers, and the resulting tokens were concatenated and paired with text tokens as input to the language model. We excluded yes/no questions from M3D-VQA to reduce binary-response bias. This stage used a batch size of 16, learning rate of $1\times 10^{-4}$, no weight decay, and was trained for 3 epochs.
Finally, during VLM fine-tuning, we used LoRA [13] with a rank of 16, scaling factor $\alpha=32$, and dropout of 0.05, updating only the LoRA modules and the projector while keeping all other weights frozen. We fine-tuned the model on both M3D-Cap and M3D-VQA (including yes/no questions), using a batch size of 8, learning rate of $5\times 10^{-5}$, no weight decay, and training for 5 epochs. All M3D-LaMed results are reported as in the original paper [2].
@@ 4.2.1 Quantitative Performance Evaluation
Interestingly, although Med3DVLM significantly outperforms RadFM and M3D-LaMed on n-gram-based metrics, all three methods achieve comparable BERTScores. This suggests that while the generated texts may be similarly aligned with reference reports at a high semantic level, Med3DVLM produces outputs with more precise lexical choices, better grammar, and clearer structure. The discrepancy between BERTScore and traditional metrics highlights the importance of evaluating both semantic fidelity and linguistic quality—especially in medical reporting, where accuracy and clinical readability are essential.
@@ 4.3.1 Quantitative Performance Evaluation
Med3DVLM achieves state-of-the-art performance in VQA, as shown in Tables 3 and 4. It consistently outperforms all baselines across both open-ended and closed-ended tasks, demonstrating superior capability in interpreting and reasoning over 3D medical images. The model achieves particularly strong results in categories such as organ and abnormality identification, reflecting its ability to extract detailed anatomical and contextual information.
Med3DVLM demonstrates greater improvements in open-ended VQA than in closed-ended tasks (9.5% in METEOR vs. 5.2% in accuracy), highlighting its strength in generating clinically relevant free-form responses that require a deeper understanding of medical imaging and context. Its strong METEOR performance suggests that its responses are not only well-aligned with ground truth but also more coherent and fluently structured than those from RadFM and M3D-LaMed. In contrast, the smaller performance gap in closed-ended VQA suggests that classification-based answering relies more on direct pattern recognition than on the model’s ability to generate expressive and context-aware responses.

### v2

@@ Abstract
Vision-language models (VLMs) have shown promise in 2D medical image analysis, but extending them to 3D remains challenging due to the high computational demands of volumetric data and the difficulty of aligning 3D spatial features with clinical text. We present Med3DVLM, a 3D VLM designed to address these challenges through three key innovations: (1) DCFormer, an efficient encoder that uses decomposed 3D convolutions to capture fine-grained spatial features at scale; (2) SigLIP, a contrastive learning strategy with pairwise sigmoid loss that improves image-text alignment without relying on large negative batches; and (3) a dual-stream MLP-Mixer projector that fuses low- and high-level image features with text embeddings for richer multi-modal representations. We evaluated our model on the M3D dataset, which includes radiology reports and VQA data for 120,084 3D medical images. The results show that Med3DVLM achieves superior performance on multiple benchmarks. For image-text retrieval, it reaches 61.00% R@1 on 2,000 samples, significantly outperforming the current state-of-the-art M3D-LaMed model (19.10%). For report generation, it achieves a METEOR score of 36.42% (vs. 14.38%). In open-ended visual question answering (VQA), it scores 36.76% METEOR (vs. 33.58%), and in closed-ended VQA, it achieves 79.95% accuracy (vs. 75.78%). These results demonstrate Med3DVLM’s ability to bridge the gap between 3D imaging and language, enabling scalable, multi-task reasoning across clinical applications. Our code is publicly available at https://github.com/mirthAI/Med3DVLM.
@@ 1 Introduction
Medical image analysis plays a crucial role in diagnosing and treating diseases such as cancer, cardiovascular conditions, and neurological disorders. However, existing models are often task-specific, lack adaptability to new tasks, and do not support real-time user interactions. Vision-language models (VLMs), such as Contrastive Language-Image Pre-Training (CLIP) [28] and Large Language and Vision Assistant (LLaVA) [20], offer greater versatility in dynamic clinical settings by aligning medical images with textual reports. CLIP leverages contrastive learning on image-text pairs to enable zero-shot image classification and image-text retrieval. LLaVA extends CLIP by integrating its visual encoder with a large language model (LLM), facilitating interactive tasks such as report generation and visual question answering (VQA). Despite their success in analyzing 2D medical images, such as chest X-rays [8] and pathology slides [22], their application to 3D imaging remains limited.
@@ 2.1 Medical Vision-Language Models
Recent efforts have aimed to develop VLMs for 3D medical imaging. RadFM [39] introduced a generalist VLM trained on massive datasets that consisted of both 2D and 3D medical images and associated radiology reports. It integrated contrastive learning and generative modeling across diverse imaging modalities (X-ray, CT, MRI) and textual reports, enabling a unified radiology representation. Extending this, M3D [2] introduced a large-scale dataset, M3D-Data, along with M3D-LaMed, a multi-modal vision-language model for 3D medical image analysis tasks such as image-text retrieval, report generation, VQA, and promotable segmentation. Its successor, E3D-GPT [15], employed a 3D multi-modal masked autoencoder framework to further enhance image-text retrieval, leading to improved performance in report generation and VQA. In parallel, researchers have also explored text-guided 3D medical image segmentation [48, 14, 40], where language prompts are used to localize anatomical structures or pathologies.
@@ 3.2.1 Contrastive Pretraining
Such a decomposition strategy enables the processing of 3D volumes at a larger size of 128x256×256 in this study, preserving fine-grained spatial details for improved image-text alignment. DCFormer, short for DeComposed Former, derives its name from this decomposition strategy. The “Former” suffix is inspired by the MetaFormer [43] framework, which generalizes Transformer-like architectures by replacing attention mechanisms with alternative token-mixing operations.
In this work, we utilize DCFormer-small variant, which consists of a stem stage and four hiararchical stages with [2, 3, 6, 2] layers and [96, 192, 384, 768] channels [1]. This adaptation of DCFormer within the LLaVA [20] framework enables efficient processing of high-resolution 3D volumes while preserving spatial detail. By reducing computational complexity, it supports scalable 3D medical vision-language modeling under typical GPU constraints and addresses key limitations of 3D vision transformers and CNNs.
@@
For comparison, the CLIP loss is given by:
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{2|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\left(\underbrace{\log\frac{e^{t\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j=1}^{|\mathcal{B}|}e^{t\mathbf{x}_{i}\cdot\mathbf{y}_{j}}}}_{\text{image}\rightarrow\text{text softmax}}+\underbrace{\log\frac{e^{t\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j=1}^{|\mathcal{B}|}e^{t\mathbf{x}_{j}\cdot\mathbf{y}_{i}}}}_{\text{text}\rightarrow\text{image softmax}}\right)$ (6)
Unlike traditional medical VLMs that rely on CLIP pretraining and require large batch sizes for effective contrastive learning, SigLIP adopts a pairwise sigmoid loss that operates independently on each image-text pair. Rather than comparing one positive against many negatives using softmax, SigLIP formulates contrastive learning as a binary classification problem. Matched pairs are pushed together, and unmatched pairs are pushed apart. This design removes the need for batch-wide normalization, improves alignment stability, and reduces sensitivity to batch size, which makes it more suitable for small-batch, semantically rich medical imaging settings.
@@ 3.2.2 Multi-modal Projector Pretraining
We propose an effective multi-modal projector for the interaction between visual and textual embeddings. In this pretraining stage, all model weights are frozen except for those in the projector. To effectively capture spatial and feature dependencies, we adopt the MLP-Mixer architecture [30], which alternates between a token-mixing MLP (spatial mixer) and a channel-mixing MLP.
@@
To further enhance multi-modal fusion, we propose a low-high hybrid Mixer-MLP architecture (Figure 2), inspired by Janus Pro [5]. Unlike M3D-LaMed [2], which produces large spatial outputs $(B,2048,768)$ that require lossy downsampling, DCFormer outputs semantically rich features $(B,32,768)$ from its final layer and spatially detailed features $(B,256,384)$ from its penultimate layer. Because of the compact spatial dimensions of DCFormer’s outputs, our model can adopt the low-high hybrid structure without exceeding the token length limit of the LLM. This is crucial, as excessively long image token sequences can crowd out text tokens, leading to truncated questions or reports during inference. To leverage both types of features, we utilize two parallel MLP-Mixer modules—one for each layer. After passing through $N$ Mixer layers, the resulting feature outputs are concatenated into a unified sequence of image tokens. In parallel, the input text is tokenized into a sequence of token IDs using the vocabulary of the LLM. This step follows the LLaVA [20] framework, which requires token embeddings for multimodal fusion. The resulting image and text tokens are then fused and passed to the LLM for joint reasoning.
This hybrid design enables the projector to fuse high-level abstract semantics with low-level spatial details, improving multi-modal alignment. Unlike prior approaches that downsample features at the expense of context, our method preserves diverse information without increasing token length. As a result, our projector produces richer and more semantically aligned joint representations, significantly enhancing radiology report generation and VQA. To the best of our knowledge, this work is the first to introduce a dual stream low-high hybrid MLP-Mixer for multimodal fusion in a 3D vision-language model. Unlike the original single-stream MLP-Mixer, our design processes both low-level spatial features and high-level semantic representations extracted from the 3D image encoder. This configuration enables the model to align textual inputs with both fine-grained anatomical details and high-level semantic context, which is critical for medical vision-language understanding. Empirically, it outperforms standard single-scale MLP-Mixer across multiple tasks.
@@
We conducted this pretraining stage on both the M3D-Cap and M3D-VQA datasets. To mitigate binary-response bias and ensure richer semantic grounding, we excluded yes/no questions from M3D-VQA.
@@ 3.3 Evaluation Metrics
We evaluate the quality of radiology report generation and open-ended VQA using four key metrics: BLEU [25], ROUGE [17], METEOR [3], and BERTScore [46]. BLEU measures precision by evaluating the overlap of n-grams between the generated text and the reference text, focusing on how many words or phrases in the generated report are found in the reference text. The BLEU score is calculated as:
$\text{BLEU}=BP\cdot\exp\left(\sum_{n=1}^{N}w_{n}\cdot\log p_{n}\right)$ (10)
where BP (brevity penalty) is defined as:
$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-\frac{r}{c})}&\text{if }c\leq r\end{cases}$ (11)
Here, $c$ is the length of the generated text, $r$ is the length of the reference text, $w_{n}$ is the weight for n-grams, and $p_{n}$ is the precision for n-grams of size $n$.
ROUGE emphasizes recall, assessing how much of the reference text appears in the generated output, which is particularly useful for summarization tasks to ensure key details are retained. The ROUGE score is calculated as:
$\text{ROUGE}=\frac{\sum_{n=1}^{N}\text{Recall}_{n}}{\sum_{n=1}^{N}\text{Reference}_{n}}$ (12)
where $\text{Recall}_{n}$ is the recall of n-grams in the generated report, and $\text{Reference}_{n}$ is the reference n-grams.
METEOR improves upon BLEU and ROUGE by considering synonyms and stemming, balancing precision and recall to provide a more flexible and comprehensive measure. The METEOR score is calculated as:
$\text{METEOR}=\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (13)
where Precision and Recall are calculated based on the number of matching words between the generated and reference reports.
BERTScore utilizes deep learning-based contextual embeddings to evaluate semantic similarity. The BERTScore is calculated as:
$\text{BERT-Score}=\frac{1}{N}\sum_{i=1}^{N}\frac{\sum_{j=1}^{M}\text{cosine}(e_{i},e_{j})}{M}$ (14)
where $e_{i}$ and $e_{j}$ are the contextual embeddings of the generated and reference words, respectively, and $M$ is the number of reference words.
@@ 3.4 Implementation Details
We implemented Med3DVLM using PyTorch [26] and Hugging Face Transformers [38], leveraging DeepSpeed ZeRO2 and BF16 precision on 8 NVIDIA A100 80GB GPUs. We used the AdamW optimizer [21] with a warmup ratio of 0.03 and a cosine learning rate scheduler.
3.4.1 Contrastive Pretraining
During contrastive pretraining on the M3D-Cap dataset, we used image-text pairs, where images were 3D CT volumes and texts were corresponding radiology reports. We computed similarity scores between pooled image and text embeddings to optimize a contrastive loss. We resized 3D CT volumes to $128\times 256\times 256$. We used DCFormer-small [1] as the vision encoder, and ClinicalBERT [34] as the text encoder. Image features were aggregated via mean pooling, while text features were achieved from the [CLS] token, as it captures the global semantic representation of the entire sequence in ClinicalBERT, making it suitable for contrastive alignment with image features. Both were projected to a shared 768-dimensional space. We used a batch size of 64, learning rate of $1\times 10^{-4}$, weight decay of 0.1, and trained for 100 epochs.
3.4.2 VLM Pretraining
For VLM pretraining, we used image-question-answer triplets, where the image was 3D CT volume, the question was a natural language question query, and the answer was the corresponding free-text response. These triplets were sourced from the M3D-Cap and M3D-VQA datasets (excluding yes/no questions) dataset. The DCFormer-small encoder was kept frozen. To align the image encoder and the LLM, we introduced a hybrid MLP-Mixer projector that fused low-level features from the penultimate encoder layer ($256\times 384$) and high-level features from the final layer ($32\times 768$) of the image encoder. Each feature stream passed through $N$ Mixer layers, and the resulting tokens were concatenated and paired with text tokens as input to the LLM. This stage used a batch size of 16, learning rate of $1\times 10^{-4}$, no weight decay, and was trained for 3 epochs.
3.4.3 VLM Fine-tuning
For VLM fine-tuning, We continued training using image–question–answer triplets from M3D-Cap and M3D-VQA, now including yes/no questions. The image encoder and MLP-Mixer projector were reused. The model was fine-tuned to generate answers conditioned on the image and question. We used LoRA [13] with a rank of 16, scaling factor $\alpha=32$, and dropout of 0.05, updating only the LoRA modules and the projector while keeping all other weights frozen. We fine-tuned the model using a batch size of 8, learning rate of $5\times 10^{-5}$, no weight decay, and training for 5 epochs. All M3D-LaMed results are reported as in the original paper [2].
3.4.4 Inference and Evaluation
During inference, we evaluated two tasks: report generation and visual question answering. In both tasks, the input consisted of a 3D CT volume and a natural language prompt. For report generation, the prompt was a fixed instruction such as “Please provide a radiology report for the given CT volume.” The model generated a free-text radiology report (see Fig 3 for an example). In VQA, the prompt was a question about the image. In open-ended VQA, the model generated free-form answers. In close-ended VQA, multiple answer options were included in the prompt, and the model selected the most appropriate ones (see Fig 4 for examples).
@@ 4.2.1 Quantitative Performance Evaluation
Interestingly, although Med3DVLM and M3D-LaMed achieve similar BERTScores (88.11 vs. 88.46), BERTScore primarily reflects abstract similarity in latent space and does not fully capture linguistic precision or clinical completeness. In contrast, Med3DVLM shows substantial gains on semantically grounded metrics, with METEOR improving from 14.38% to 36.42% and ROUGE from 19.55% to 40.25%. These improvements indicate that Med3DVLM generates reports with more accurate lexical choices, clearer structure, and better alignment with clinical content. The discrepancy between BERTScore and traditional n-gram metrics highlights the importance of evaluating both semantic fidelity and textual quality, which are essential for accurate and readable medical reporting.
@@ 4.3.1 Quantitative Performance Evaluation
Med3DVLM achieves state-of-the-art performance in both open-ended and close-ended VQA tasks, as shown in Tables 3 and 4. The model demonstrates particularly strong results in clinically important categories such as organ and abnormality identification, highlighting its ability to capture fine-grained anatomical and contextual cues in 3D volumes.
In open-ended VQA, Med3DVLM improves the METEOR score from 33.58% to 36.76% (a 9.5% improvement) and the ROUGE score from 52.39% to 56.31% (a 7.5% improvement) compared to M3D-LaMed. These metrics offer a more semantically grounded evaluation than simple lexical overlap, capturing both the accuracy and fluency of generated responses, which are essential for answering complex clinical queries with clarity and relevance. The results suggest that Med3DVLM produces free-text responses that are not only better aligned with the ground truth but also more structured and clinically coherent.
In close-ended VQA, Med3DVLM achieves 79.75% accuracy, surpassing M3D-LaMed’s 75.78% by 5.2%. Although the margin is smaller than in open-ended VQA, the improvement remains clinically meaningful: even modest gains in accuracy can reduce diagnostic errors and enhance the reliability of automated decision support systems in clinical practice.
@@ 4.4 Ablation Study
4.4.1 Effect of Core Components
Table 5: Effect of core Med3DVLM components on radiology report generation.
DCFormer SigLIP Loss MLP-Mixer BLEU ROUGE METEOR BERT
24.47 28.80 23.78 85.75
14.31 18.14 13.48 83.86
8.62 11.23 8.88 77.07
36.88 40.25 36.42 88.11
Table 6: Effect of core Med3DVLM components on open-ended VQA performance.
DCFormer SigLIP Loss MLP-Mixer Metric Plane Phase Organ Abnormality Location Mean
BLEU 99.00 77.28 39.13 19.73 24.85 52.00
ROUGE 99.05 84.23 43.48 23.85 29.48 56.02
METEOR 49.52 68.31 27.23 16.69 20.07 36.36
BERT 99.85 96.30 90.33 86.29 87.74 92.10
BLEU 98.60 75.00 34.34 17.76 24.55 50.05
ROUGE 98.67 81.55 38.59 21.45 28.84 53.82
METEOR 49.37 65.18 24.28 14.72 19.69 34.65
BERT 99.79 95.87 89.44 85.93 87.71 91.75
BLEU 0.00 0.17 0.19 0.73 0.57 0.33
ROUGE 0.08 0.43 0.96 1.54 1.46 0.89
METEOR 0.07 0.90 1.62 2.19 2.14 1.38
BERT 73.18 74.68 74.47 74.01 74.24 74.12
BLEU 98.85 78.17 40.22 18.99 25.66 52.38
ROUGE 98.89 84.20 45.22 23.27 29.99 56.31
METEOR 49.43 68.50 29.32 16.21 20.32 36.76
BERT 99.83 96.47 90.47 86.27 87.88 92.18
Table 7: Effect of core Med3DVLM components on closed-ended VQA accuracy.
DCFormer SigLIP MLP-Mixer Plane Phase Organ Abnormality Location Mean
99.30 87.95 77.40 69.90 63.99 79.71
98.85 84.95 74.80 69.82 64.46 78.58
1.60 1.05 1.35 1.03 1.28 1.26
99.15 87.50 77.45 70.17 64.49 79.75
We first analyze the impact of the three core components of Med3DVLM: the DCFormer vision encoder, the SigLIP contrastive loss, and the multi-scale MLP-Mixer projector. The results are summarized in Table 5, Table 6, and Table 7, corresponding to radiology report generation, open-ended VQA, and closed-ended VQA, respectively.
In the radiology report generation task, the full model achieves a METEOR score of 36.42%. Removing the MLP-Mixer leads to a drop in performance to 23.78%, while removing the SigLIP loss results in a lower METEOR score of 13.48%. When the DCFormer encoder is removed, the MLP-Mixer must also be excluded since it depends on the multi-scale feature outputs from DCFormer. In this case, the METEOR score degrades severely, reaching only 8.88%. Similar trends are observed in BLEU, ROUGE, and BERTScore, highlighting the critical role of all three components in producing clinically accurate and semantically aligned text.
For open-ended VQA, removing both DCFormer and MLP-Mixer leads to a drastic drop in performance, reducing the mean METEOR score from 36.76% to just 1.38%. In contrast, removing only the MLP-Mixer causes a modest drop to 34.65%, while removing SigLIP results in a minimal decrease to 36.36%. Similarly, for closed-ended VQA, the full model achieves a mean accuracy of 79.75%. Removing SigLIP has little effect (79.71%), and removing the MLP-Mixer causes a small drop to 78.58%. In contrast, removing both DCFormer and MLP-Mixer results in a sharp decline to 1.26% accuracy. These results suggest that while all three components contribute to overall VQA performance, the DCFormer encoder is essential for reliable visual understanding, and the MLP-Mixer plays an important role in multi-modal reasoning.
In summary, each core component of Med3DVLM contributes to overall performance, with varying importance across tasks. The DCFormer encoder is essential for robust 3D representation across all settings, while the MLP-Mixer plays a key role in multi-modal reasoning, particularly for generative tasks. The SigLIP loss enhances alignment stability, especially in report generation.